11. The bottom of a 20' ladder is placed 5' away from a structure. Where will the top of the ladder touch the building?
Answer by stanbon(75887)

The bottom of a 20' ladder is placed 5' away from a structure. Where will the top of the ladder touch the building?
----------
Draw the picture.
You should have a right triangle with the right angle where the ground
meets the building.
The hypotenuse is 20'
One of the legs is 5'
Let the other leg be "x"
EQUATION:
x^2 + 5^2 = 20^2
x ^2 = 400 - 25
x^2 = 375
x = sqrt(375) = sqrt(25*15) = 5sqrt15 ft.
==============
Cheers,
Stan H.